### v3

# Mercury: A Code Efficiency Benchmark for Code Large Language Models

Abstract
Amidst the recent strides in evaluating Large Language Models for Code (Code LLMs), existing benchmarks have mainly focused on the functional correctness of generated code, neglecting the importance of their computational efficiency. To fill the gap, we present Mercury, the first code efficiency benchmark for Code LLMs. It comprises 1,889 Python tasks, each accompanied by adequate solutions that serve as real-world efficiency baselines, enabling a comprehensive analysis of the runtime distribution. Based on the distribution, we introduce a new metric Beyond, which computes a runtime-percentile-weighted Pass score to reflect functional correctness and code efficiency simultaneously. On Mercury, leading Code LLMs can achieve 65% on Pass, while less than 50% on Beyond. Given that an ideal Beyond score would be aligned with the Pass score, it indicates that while Code LLMs exhibit impressive capabilities in generating functionally correct code, there remains a notable gap in their efficiency. Finally, our empirical experiments reveal that Direct Preference Optimization (DPO) serves as a robust baseline for enhancing code efficiency compared with Supervised Fine Tuning (SFT), which paves a promising avenue for future exploration of efficient code generation.  [^1] Our code and data are available on GitHub: https://github.com/Elfsong/Mercury.
1 Introduction
The domain of code generation, which aims to empower computers to autonomously generate code based on natural language task descriptions (NL2Code), has long been considered a promising way to facilitate interaction between humans and computers [46, 44]. The recent emergence of Large Language Models (LLMs) has spurred a new wave of NL2Code models [34, 38, 12, 29, 4], which leverage the impressive language understanding and generative capabilities of LLMs to drive forward the ambitious goal of synthesizing high-quality code from natural language instructions.
To measure the quality of code, recent code generation benchmarks mainly focus on evaluating their functional correctness via test case fuzzing [27]. This approach assesses the outcome congruence between the LLM-generated and canonical solutions by executing bespoken test cases. For instance, HumanEval [8] and MBPP [3] collected a small but fine set of handcrafted tasks with test cases. EvalPlus [26] further consolidates these two above benchmarks by augmenting the case scope. On the contrary, APPS [15] widely gathered over 5,000 public coding tasks from online platforms. Despite these strides, there is a discernible oversight in current code generation benchmarks concerning the code efficiency evaluation, although that is critical in software development [45, 47]. Moreover, handcrafting diverse solutions and test cases to cover all scenarios is infeasible [26]. In light of these findings, we highlight vital limitations inherent in the existing code generation benchmarks:
Figure 1: Executing these two LLM-generated codes on 100 test cases. While both codes successfully follow the task instruction and pass all test cases, the right snippet notably excels in code efficiency, completing in a mere 121 ms compared to the 5,714 ms consumed by the left snippet. As Code LLMs become widely used in the real world, code efficiency determines factual productivity, where Mercury can gauge the vital metric.
  1. 1.
    Absence of code Efficiency Evaluation. Existing code generation benchmarks focus on assessing functional correctness while overlooking the evaluation of code efficiency [8, 3, 15]. As illustrated in Figure 1, despite both code snippets can handle the sorting task functionally, the right efficient solution (121 ms) is nearly 50 times faster than the left inefficient solution (5,714 ms). This striking runtime differentiation underscores the necessity of incorporating code efficiency assessments within code generation benchmarks, encouraging Code LLMs to produce not only correct but also efficient code.
  2. 2.
    Insufficient Test Case Coverage. As shown in Table 1, most code generation benchmarks manually build a small number of test cases or extract the accompanying test cases from existing resources, potentially overlooking edge cases and nuanced code behaviors [8, 3]. For example, Figure 8 displays that HumanEval #55 contains only 3 test cases, testing up to the 12th Fibonacci number [8]. Its given canonical solution will quickly reach the recursion depth limitation when computing a larger Fibonacci number (the recursion limitation depends on the environment). Therefore, notwithstanding the generated code satisfies all test cases, such success does not necessarily equate to assurance of functional correctness and much less to code efficiency.
  3. 3.
    Lack of Task Diversity. Another noticeable deficit of existing code generation benchmarks is the insufficient diversity and complexity in their tasks [8, 3, 27]. Since most benchmarks only consist of elementary-level programming tasks, recent Code LLMs can effortlessly tackle most tasks regardless of their actual capacities [47]. This flaw results in these benchmarks failing to pose a substantial challenge to Code LLMs and truly reflect their underlying potential.
Code Efficiency. Code efficiency refers to the performance measure of time and space complexity to accomplish a specific task. Efficient code can improve user experience, save energy, and make applications more sustainable and cost-effective. Compared with the scalable memory space, execution time is the performance bottleneck of most codes. Consequently, this work focuses on the time dimension of code efficiency.
Figure 2: An overview of Mercury dataset. Each Mercury task has a task description, a test case generator, a prompt & entry point, and corresponding solutions. To evaluate code efficiency, we introduce the Beyond metric, which signifies the runtime percentile of the LLM-generated code on the runtime distribution supported by corresponding solutions. In this example, the LLM-generated code executes in 521 ms, outpacing 86.18% of collected solutions on the runtime distribution. Consequently, the Beyond metric in this case is 86.18%.
Our Benchmark. In this work, we introduce Mercury, a novel code generation benchmark designed to assess and improve the code efficiency of Code LLMs. As depicted in Figure 2, Mercury comprises 1,889 Python programming tasks with three difficulty stratification, which is divided into two datasets for model evaluation and fine-tuning separately. For each evaluation task, we assign a test case generator to remedy the shortfall of test case coverage. In measuring code efficiency, the primary challenge stems from normalizing the absolute runtime across tasks that have diverse runtime ranges. Thus, we collect and locally execute numerous historical solutions for each task to form a runtime distribution and leverage the runtime percentile of LLM-generated code on the distribution instead of the absolute runtime to evaluate code efficiency. Furthermore, to mitigate performance discrepancies attributed to irrelevant processes and diverse hardware configurations, we set up an isolated sandbox environment for task execution to establish local runtime distributions.
Contribution. Our work aimed to fill the code efficiency evaluation gap in code generation benchmarks with the following key contributions:
  • Dataset. We collect a novel code generation dataset Mercury designed to assess and improve Code LLM code efficiency in Section 2, accompanied by an extensible open-source data collection framework for enriching Mercury with more tasks and programming languages.
  • Metric. We propose the first efficiency-focused code generation metric Beyond and establish a benchmark to evaluate leading Code LLMs using this metric in Section 3.
  • Baselines. In Section 4, we detail our extensive analysis of two baselines to enhance code efficiency while maintaining functional correctness. Experiment results reveal that despite Code LLMs excelling in functional correctness, there is still considerable potential to elevate efficiency.
Table 1: A comparison of Mercury to existing NL2Code benchmarks. Mercury distinguishes itself by including a set of distilled high-quality solutions and a dedicated test case generator for each task. * signifies that the solution number can be further expanded by the data collection framework.
Benchmarks Tasks Sources Cases Solutions Difficulty
HumanEval 164 Crowd Source 8.08 1 1
MBPP 257 Crowd Source 3.01 1 1
APPS 5,000 Online 21.2 23.4 3
Mercury 256 Online + Filters $\infty$ 18.4 * 3
2 Mercury Datasets
We initiate this work by collecting public programming tasks on Leetcode [23]. Subjecting these questions to a series of filters, we distilled them down to 1,889 high-quality tasks. A difficulty-balanced subset of 256 tasks was randomly selected to form the Mercury-eval benchmark, which obtains an average of 18.4 solutions for each problem. The remaining tasks have been designated as the Mercury-train dataset for baseline training (detailed data distribution is listed in Appendix Table 6). To enhance clarity within this paper, we employ Mercury to denote Mercury-eval unless otherwise specified.
Data Schema. As illustrated in Figure 2, Mercury offers a unified data schema to streamline the evaluation procedure and bolster further development endeavors. The data scheme encompasses these principal components: (1) Task Description contains the task instruction interpreted into a plain text format, along with illustrative examples and constraints of inputs and outputs. (2) Test Case Generator refers to a Python code snippet designed to automatically produce a comprehensive set of test cases in accordance with the specifications laid out in the task description. (3) Solutions are sampled from Leetcode historical submissions. Each solution within Mercury has undergone rigorous testing, and Locality-Sensitive Hashing [17] is employed to prevent the inclusion of any identical solutions. (4) Prompts and Entry Points where prompts act as the initiating prefixes for LLM code generation and entry points denote the start point for code execution. We delineate the definition of Mercury fields in the Appendix Table 5.
Task Filters. Mercury tasks originate from public programming problems on Leetcode. To assure the quality and uniformity of the dataset, we distilled gathered tasks based on the following conditions:
  1. 1.
    Number of Solutions. To establish a solution runtime distribution for each task, we filtered out tasks having less than two associated solutions. After excluding these tasks, Mercury tasks possess an average of 18.4 unique solutions.
  2. 2.
    Restricted Data Structure. Above the inherent Python data types, Mercury also incorporates two custom data types: Binary Tree and Linked List (the specific structure definitions can be found in Appendix Figure 4), which increases Mercury’s diversity and escalates its difficulty level. Tasks that contain other data structures will be removed.
  3. 3.
    Unique Outputs. Certain Leetcode tasks may permit non-unique answers. For example, a result list can be returned in any order. Evaluating all possible answers can drastically complicate the test case verification process. To eliminate this problem, we harness the corresponding test case generator to generate $N$ test cases $T_{i}=\langle Input_{i},Output_{i}\rangle\ s.t.\ i\in\{0,1,\cdots,N\}$ and execute $T$ on different solutions $S_{m}\ s.t.\ m\in\{0,1,\cdots,M\}$ to observe if all $Output_{i}=S_{m}(Input_{i})\ s.t.\ i\in\{0,1,\cdots,N\}$ remain identical. Any tasks that potentially yield non-unique answers were subsequently excluded.
Task Difficulty. Most existing NL2Code benchmarks predominantly comprise simplistic tasks, leading to a situation where LLMs of varied capabilities address most tasks effortlessly and yield indistinguishable high scores [47, 18]. To alleviate this issue, Mercury inherits the difficulty categorization from Leetcode, i.e., Easy, Medium, and Hard. The stratification aims to probe the upper bounds of Code LLM capabilities, delivering a more evident distinction between various Code LLMs.
Test Case Generator. Manual creation of test cases can be a laborious process. To gather sufficient test cases to conduct an exhaustive assessment, we assign a test case generator for each evaluation task, which can produce a full range of test cases to thoroughly evaluate the functional correctness and code efficiency of given solutions. Specifically, We feed $pretty\_content$ into GPT-4 [34] to generate an initial test case generator snippet. To confirm the effectiveness of the initial generator, we subsequently create 24 test cases by the generator and submit these cases to the Leetcode Online Judge (OJ) system. Should any of the generated test cases not pass the LeetCode OJ validation, we manually revise the generator until all generated cases can be successfully validated.
3 Code Efficiency Metric
In the domain of software development, code efficiency can be defined as the absolute code runtime for executing a given test case set [7]. Nonetheless, a primary obstacle in benchmarking code efficiency is normalizing runtime measurements across disparate environments. For instance, a sub-optimal solution might have a faster absolute runtime on high-performance hardware than an optimal solution on low-performance hardware. Moreover, different operation systems and code interpreters may also fluctuate the code runtime. Therefore, absolute runtime fails as a consistent and reliable code efficiency benchmark metric. To address this issue, an intuitive approach involves modeling a devoted runtime distribution for each task and calculating the average runtime percentiles of LLM solution samples over the runtime distribution. With this idea in mind, we proposed a normalized code efficiency metric Beyond:
$p_{k}^{n}=\frac{max(R^{n})-clip(r_{k}^{n},min(R^{n}),max(R^{n}))}{max(R^{n})-min(R^{n})},\qquad Beyond=\frac{\sum\nolimits_{N,K}^{n=0,k=0}p_{k}^{n}}{N\cdot K}.$ (1)
Where $N$ is the total number of tasks, and $K$ denotes the size of LLM solution samples. For a specific task $n\in N$, $R^{n}$ is the runtime array corresponding to the collected historical solutions, and $r_{k}^{n}\ s.t.\ k\in K$ denotes the runtime for the $k$-th LLM solution. $clip$ is a function to constraint the value $r_{k}^{n}$ in the range $[min(R^{n}),max(R^{n})]$. Runtime is defined as the period from the solution instantiation to the evaluation across all test cases, culminating with a successful termination (More engineering details can be found in Appendix Section A.3). Since any case failure of the $k$-th solution results in $r_{k}^{n}\rightarrow+\infty$ and then $p_{k}^{n}=0$, Beyond can reflect functional correctness as well.
Untrusted Code Execution. Since most Code LLMs are trained on an extensive code corpus from unverified sources, there is an intrinsic risk that these models may produce malicious code when driven by specific meticulous prompts [8]. The direct execution of synthesized code raises significant security concerns. To alleviate the risk of running untrusted code, we engage a robust sandbox to execute code in an isolated environment. Sandbox details are deliberated in Appendix A.3
Environment-agnostic Evaluation. To ensure fair comparison across diverse configurations, we run each task $n$ with corresponding test cases locally and aggregate their runtimes into the runtime array $R^{n}$. Appendix Figure 10 illustrates the Beyond score of two LLMs (‘deepseek-coder-33b’ and ‘deepseek-coder-6.7b’) over three distinct hardware specifications: the micro-tier (0.25 CPU cores), the small-tier (0.5 CPU cores), and the standard-tier (1 CPU core). The results demonstrate that Beyond remains consistent over different hardware configurations.
4 Experiments
In this section, we present a series of baseline experiments to improve code efficiency by training on Mercury-train dataset and assessing on the Mercury-eval dataset. Our empirical study encompasses 10 open-source LLMs with a broad parameter spectrum from 1.3 to 34 billion. For each LLM, we compare the performance of the original model and two optimization strategies, Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), for their potential to optimize LLM generating functionally correct and computationally efficient code. Finally, we analyzed the underlying factors contributing to the failure of LLMs on the Mercury-eval dataset.
4.1 Baselines
Supervised Fine-Tunning. Within the SFT [6] method, an LLM undergoes additional training on a small dataset, which aims to specialize the LLM to perform better on certain tasks correlated to the training dataset. To optimize the code efficiency performance of Code LLMs, the most intuitive strategy is to fine-tune the Code LLM using optimal runtime solutions. In our experimental setup, we apply a unified prompt template for each Code LLM to ensure a fair comparison. The “pretty_content” attribute fills the <task_content> placeholder, the “prompt” attribute fills the <code_starter> placeholder, and the <code_completion> placeholder is completed with the fastest solutions. To steer Code LLMs towards generating the intended code completion format, we prepend a one-shot example to the prompt template. Appendix Figure 9 presents the prompt template.
Direct Preference Optimization. Although SFT exemplifies a straightforward approach, it is susceptible to the pitfall of catastrophic forgetting [20]. To enable LLMs to align with human preferences while preserving their functional capabilities, existing methodologies employ reinforcement learning with human preference feedback (RLHF). However, RLHF introduces additional model complexities and potential instabilities, necessitating significant computing resources and extra reward model training [48, 5, 40]. DPO [36] bypasses these challenges by explicitly mapping reward functions and the optimal objective. This connection demonstrates that maximizing rewards under specific constraints can be effectively addressed through a singular training phase based on data reflecting human preferences. The DPO training procedure is elaborated in Appendix Section A.4.
4.2 Functional Correctness Benchmarks
HumanEval assesses the functional correctness of synthesized code derived from docstrings. It contains 164 distinct Python tasks that cover several programming areas, such as language comprehension, algorithm development, and simple mathematics [8]. MBPP has a sanitized collection of 257 entry-level Python programming problems. Each problem in this dataset consists of three components: a task description, an associated code solution, and three automated test cases to validate the code functionality [3]. Both HumanEval and MBPP harness the metric Pass to measure the Code LLMs’ functional correctness, where a task is considered solved if the given solution passes all test cases, and the total fraction of solved tasks is reported as $Pass=N_{solved}/N_{total}$ [21].
4.3 Experimental Setups
[…]
Training Data. We use Mercury-train for model training. As for the SFT process, we nominate the fastest solution as the supervised label, then format the training data as $\langle pretty\_content,prompt,solution\_optimal\rangle$. Regarding the DPO procedure, we select the top 5 pairs of solutions that exhibit the most significant discrepancy in runtime. The training date format is $\langle pretty\_content,prompt,solution\_fast,solution\_slow\rangle$.
4.4 Empirical Results
Functional correctness is the prerequisite for evaluating code efficiency for the code generation task. Our primary objective is to enhance the code efficiency without compromising the functional correctness. To this end, we first introduce the existing metric Pass to gauge the functional correctness [8] and then leverage Beyond to provide a holistic evaluation, encompassing both code efficiency and functional correctness. Finally, we measure the Gap between Beyond and Pass to mirror the baseline ability for improving efficiency while preserving correctness. These experiments aim to investigate the innate capabilities of cutting-edge Code LLMs and their potential after baseline fine-tuning. Therefore, extensive parameter optimization and prompt engineering were not pursued. To deliver a comprehensive evaluation, we have further integrated the HumanEval and MBPP benchmarks as supplementary measures for appraising functional correctness [8, 3].
Figure 3: The horizontal axis represents the score for functional correctness, while the vertical axis indicates the score for code efficiency. The left figure illustrates the performance of the baseline model, whereas the right one depicts the performance after DPO tuning. Model points located nearer to the diagonal line exhibit a more equitable balance between functionality and efficiency.
Functional Correctness. Table 2 lists Pass scores over various Code LLMs, showing that larger models tend to provide better functional correctness. Except for the smallest model “deepseek-coder-1.3b-base”, DPO invariably enhances the overall Pass scores across most Code LLMs, while SFT diminishes functional correctness on the largest two Code LLMs. These findings suggest that smaller models may struggle to integrate new knowledge while preserving their original functionality, and SFT may induce catastrophic forgetting in the pursuit of heightened code efficiency. Moreover, it is evident on Mercury that Pass scores of each model consistently decline as the difficulty level increases, indicating that the Mercury difficulty stratification is effective at probing the upper limitation of each Code LLM compared to the auxiliary benchmarks.
Table 2: Functional correctness (Pass) evaluation results. The underlined values denote the top-performed approaches among the original model and baselines. The bolded values denote the best performance on each benchmark. We sample one solution for each task to calculate pass score.
Mercury
Model Name HumanEval MBPP Easy Medium Hard Overall
deepseek-coder-1.3b-base 28.7 55.4 60.7 52.8 23.2 38.1
+ SFT 24.2 46.2 58.9 53.6 25.3 36.2 (-1.9)
+ DPO 29.1 50.2 61.4 53.6 20.0 35.9 (-2.2)
starcoder2-3b 31.7 57.4 56.1 52.1 21.6 37.8
+ SFT 29.0 47.2 60.7 58.8 25.3 38.8 (+1.0)
+ DPO 33.5 59.6 62.5 61.0 23.4 41.1 (+3.3)
deepseek-coder-6.7b-base 47.6 70.2 69.3 68.9 56.1 61.0
+ SFT 56.1 59.6 69.1 71.4 57.7 62.2 (+1.2)
+ DPO 54.3 72.8 74.1 72.6 58.9 65.4 (+4.4)
starcoder2-7b 35.2 54.4 63.6 61.7 29.2 44.3
+ SFT 42.9 57.2 64.8 58.5 31.3 47.5 (+3.2)
+ DPO 55.4 61.4 74.8 66.9 32.6 53.6 (+9.3)
CodeLlama-7b-hf 33.5 52.0 55.7 41.7 12.9 29.6
+ SFT 29.5 47.6 58.9 38.5 16.1 31.3 (+1.7)
+ DPO 38.7 49.2 67.5 45.7 17.9 36.1 (+6.5)
CodeQwen1.5-7B 51.8 72.2 70.0 70.1 49.7 61.1
+ SFT 54.3 74.8 70.9 67.9 49.7 61.9 (+0.8)
+ DPO 55.5 75.4 72.5 66.9 45.7 61.1 (+0)
starcoder2-15b 46.3 66.2 69.5 65.4 50.3 58.0
+ SFT 51.6 69.2 72.0 68.9 51.7 61.3 (+3.3)
+ DPO 57.0 72.8 78.0 73.8 54.7 66.7 (+8.7
CodeLlama-13b-hf 37.8 62.4 76.8 60.5 18.4 39.6
+ SFT 39.5 59.8 65.5 54.8 19.5 39.5 (-0.1)
+ DPO 49.1 64.4 78.6 60.0 29.0 50.1 (+10.6)
deepseek-coder-33b-base 54.3 73.2 70.9 67.9 62.3 65.0
+ SFT 58.1 74.8 61.8 58.0 47.1 58.7 (-6.3)
+ DPO 72.9 80.6 78.9 76.5 61.6 73.4 (+8.4)
CodeLlama-34b-hf 48.2 65.4 77.7 63.7 32.4 52.4
+ SFT 52.8 68.2 61.8 58.0 26.2 47.5 (-4.9)
+ DPO 65.9 75.2 83.9 68.4 63.2 70.6 (+18.2)
Table 3: Code efficiency (Beyond) evaluation results across three difficulty levels. The bolded value indicates the top performance for each metric, while the underlined values denote the most effective approaches among the original model and the baselines. In our experiment, we sample 5 solutions for each task to calculate Beyond score.
Model name Easy Medium Hard Overall Gap
deepseek-coder-1.3b-base 47.97 39.77 19.26 35.62 9.85
+ SFT 42.58 38.12 18.67 33.04 (-2.58) 12.74 (+2.89)
+ DPO 46.91 42.27 16.78 35.21 (-0.41) 9.64 (-0.21)
starcoder2-3b 43.55 41.91 15.21 33.40 9.72
+ SFT 44.64 42.10 15.72 34.01 (+0.61) 14.04 (+4.31)
+ DPO 43.70 41.02 12.99 32.42 (-0.99) 16.33 (+6.61)
deepseek-coder-6.7b-base 48.80 51.16 45.11 48.29 16.40
+ SFT 51.37 52.71 44.28 49.39 (+1.09) 16.55 (+0.16)
+ DPO 56.25 52.35 40.62 49.70 (+1.41) 18.73 (+2.34)
starcoder2-7b 50.23 51.29 20.25 40.37 10.95
+ SFT 42.21 44.02 21.09 35.61 (-4.77) 15.80 (+4.84)
+ DPO 53.52 51.41 17.35 40.56 (+0.18) 17.41 (+6.46)
CodeLlama-7b-hf 42.55 30.99 8.88 27.45 9.27
+ SFT 39.75 26.89 9.55 25.41 (-2.04) 12.48 (+3.21)
+ DPO 54.14 34.48 11.10 33.29 (+5.84) 10.46 (1.19)
CodeQwen1.5-7B 51.11 53.56 39.03 47.78 15.35
+ SFT 54.16 51.43 38.05 47.82 (0.04) 14.91 (-0.43)
+ DPO 56.07 51.55 38.05 48.52 (0.74) 13.12 (-2.22)
starcoder2-15b 58.18 52.09 37.34 49.17 12.55
+ SFT 53.54 52.77 37.73 47.92 (-1.25) 16.22 (+3.67)
+ DPO 68.29 59.54 48.97 58.95 (9.78) 10.81 (-1.74)
CodeLlama-13b-hf 57.00 44.25 12.99 38.01 13.79
+ SFT 44.95 39.96 13.55 32.70 (-5.31) 13.78 (-0.01)
+ DPO 67.09 55.72 19.72 47.39 (9.38) 8.47 (-5.32)
deepseek-coder-33b-base 51.26 48.90 45.43 48.53 18.50
+ SFT 40.33 37.75 36.82 38.32 (-10.21) 17.30 (-1.20)
+ DPO 74.59 68.91 55.98 66.47 (+17.94) 5.79 (-12.70)
CodeLlama-34b-hf 56.28 48.21 22.96 42.40 15.49
+ SFT 45.49 44.96 20.73 36.91 (-5.50) 11.61 (-3.88)
+ DPO 78.55 60.95 51.94 63.94 (+21.54) 8.01 (-7.47)
Code Efficiency. Regarding the NL2Code task, once functional correctness has been assured, attention naturally pivots to enhancing code efficiency. As depicted in Table 3, we investigate code efficiency metrics across a spectrum of Code LLMs. Experiments demonstrate that DPO yields a stable enhancement in code efficiency from models exceeding 6.7B parameters. In contrast, SFT detracts most Beyond scores from original models, suggesting that the plain SFT may not be a feasible strategy for enhancing code efficiency. Further analysis compares the Gap between Beyond and $Pass$. Since the ideal Beyond should be aligned with Pass (where the LLM-generated solution is correct and faster than all historical solutions), it shows how much the baseline method shrinks the gap between functional correctness and code efficiency. Our findings indicate that DPO substantially narrows Gap in models larger than 15B parameters. However, Gap tends to widen in smaller models under the same configuration. This implies that larger models possess a greater capacity to assimilate the nuanced knowledge to make strides in code efficiency while retaining their functional correctness.
4.5 Failure Analysis
Table 4 provides an error breakdown of where Code LLMs misstep during the Mercury evaluation:
(1) Generation Errors arise from syntactical issues. The common manifestations include improper indentation, mismatched parentheses, or unexpected truncation. Fine-tuning introduces additional knowledge for Code LLMs to adapt the Mercury convention, emphasizing standard indentation, concise code, and minimal comments. Therefore, both SFT and DPO generally reduced these errors.
(2) Execution Errors differ from Generation Errors because they occur after the code has been successfully loaded. These errors emerge as exceptions, which could stem from various issues, such as flawed code logic, execution timeouts, memory leakage, or sandbox interruption. We observe that SFT tends to aggravate these errors on most models, whereas DPO mitigates these errors successfully.
(3) Test Case Errors are the most prevalent errors where the code is executed without exceptions, but the output fails to align with the expectation.
Table 4: The distribution of failure cases across Code Generation, Code Execution, and Test Case errors. E/M/H indicates Easy/Medium/Hard levels, respectively. We sample 5 solutions for each task, so there are $256*5=1280$ solutions in total for each model.
Model Name Code Generation Code Execution Test Case Pass
E M H E M H E M H E M H
deepseek-coder-1.3b-base 82 85 59 17 33 95 74 73 180 267 214 101
+ SFT 106 104 37 16 8 63 59 76 225 259 217 110
+ DPO 90 108 40 17 5 49 63 75 259 270 217 87
starcoder2-3b 107 102 33 35 26 107 51 66 201 247 211 94
+ SFT 97 93 24 29 13 90 47 61 211 267 238 110
+ DPO 79 75 11 30 14 87 56 69 235 275 247 102
deepseek-coder-6.7b-base 107 101 30 16 5 56 12 20 105 305 279 244
+ SFT 105 100 25 17 6 61 14 10 98 304 289 251
+ DPO 87 82 23 12 6 58 15 23 98 326 294 256
starcoder2-7b 107 101 22 21 9 74 32 45 212 280 250 127
+ SFT 105 100 22 18 13 72 32 55 205 285 237 136
+ DPO 90 90 21 10 11 61 11 33 211 329 271 142
CodeLlama-7b-hf 23 28 23 41 69 122 131 139 234 245 169 56
+ SFT 11 9 17 44 72 112 126 168 236 259 156 70
+ DPO 9 10 12 23 56 117 111 154 228 297 185 78
CodeQwen1.5-7B 105 100 18 1 4 44 26 17 157 308 284 216
+ SFT 105 101 16 4 3 35 19 26 168 312 275 216
+ DPO 98 96 26 5 8 35 18 30 175 319 271 199
starcoder2-15b 105 100 20 4 7 49 25 33 147 306 265 219
+ SFT 104 100 18 3 3 56 16 23 136 317 279 225
+ DPO 83 64 10 1 1 33 13 41 141 343 299 251
CodeLlama-13b-hf 10 14 28 19 41 99 73 105 228 338 245 80
+ SFT 46 52 32 29 19 111 77 112 207 288 222 85
+ DPO 24 9 24 10 12 100 60 141 185 346 243 126
deepseek-coder-33b-base 105 103 26 11 11 47 12 16 91 312 275 271
+ SFT 69 78 27 27 26 65 72 66 138 272 235 205
+ DPO 56 75 15 9 7 86 28 13 66 347 310 268
CodeLlama-34b-hf 22 35 50 28 55 84 48 57 160 342 258 141
+ SFT 35 97 50 37 19 56 96 54 215 272 235 114
+ DPO 4 12 10 26 76 30 41 40 120 369 277 275
5 Related Work
NL2Code Generation is the task of generating a computer program that satisfies given specifications. Initial approaches to converting natural language to code relied on rigid methods like probabilistic grammars and domain-specific languages, having limited flexibility and scalability [19, 11]. The advent of statistical models, such as n-grams and Hidden Markov models, attempted to overcome these limitations but struggled with modeling complexity and dependencies [31, 41]. The transformational impact of the Transformer model [42] and its subsequent application to NL2Code [30] led to the development of LLMs like Codex, which significantly improved the task’s feasibility by utilizing extensive unlabelled data sets [8]. Follow-up LLMs such as AlphaCode [25], CodeGen [32], PaLM-Coder [9], and StarCoder [24] continued to advance this research field, exhibiting emergent abilities in coding and debugging that mirrored human programmers.
NL2Code Evaluation currently focuses on gauging the functional correctness of generated code. As a pioneer, CodeBLEU [37] adapts the BLEU [35] metric into code generation. However, given the abstract nature of programming languages, distinct code can express the equivalent semantics, prompting subsequent benchmarks to harness test case fuzzing instead of the similarity measurement. For example, HumanEval [8] and MBPP [3] consist of hand-written Python programming tasks and corresponding test cases. On the note of enhancing language inclusiveness, ODEX [43] integrates multiple natural languages, while MBXP [2] extends the benchmarks to cater to a variety of programming languages, promoting polyglot code generation evaluation. Recent benchmarks have also begun to consider more aspects beyond functional correctness. For instance, the benchmark DS-100 [22] dives deeply into the data analysis scenarios, and CodeGen [32] contributes a benchmark for multi-turn code generation. For security-oriented code generation, SecurityEval [39] offers a concentrating benchmark on mining the vulnerability of generated code.
6 Limitations
In this work, we measure code efficiency under the assumption that the code runtime is uniformly distributed. The simplification streamlines code efficiency evaluation via limited solution samples. However, the distribution of code runtime in real-world scenarios is more intricate, which may call for more solution samples to support more precise modeling. Additionally, the presence of data contamination during the model training phase compromises the precision of the Mercury benchmark to reflect the performance of tainted models [18]. To mitigate this issue, we will update our benchmark via our open-sourced data collection framework to import new tasks dynamically, thus laying the groundwork for more detailed investigations in subsequent studies.
7 Conclusion
In this work, we introduced Mercury, the first code efficiency benchmark for NL2Code evaluation. Unlike prior work that focused on functional correctness, our benchmark highlights the importance of code efficiency. By crafting dedicated test case generators and sampling ground-truth solutions across all difficulty levels from Leetcode, we have developed a comprehensive and rigorous Code LLM evaluation frame. We evaluated leading Code LLMs against benchmarks and found that even though these models are proficient in generating functionally correct code, there is still considerable space for code efficiency improvement. As Code LLMs become more widely used, code efficiency determines factual productivity, where Mercury can gauge the vital metric. As a commitment to ongoing research and to foster further innovation in this area, we have open-sourced the Mercury dataset collection framework, laying the groundwork for future advancements in the field.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The abstract and introduction include the main contributions and the research scope.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Given that our work is largely technical and does not engage with societal systems directly, it is unlikely to have negative societal repercussions.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] The paper conforms to the ethics review guidelines.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] The paper does not include any theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] The paper does not include any theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We release our dataset on HuggingFace and our code on GitHub. See the abstract footnote.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 4.3.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Section 4.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 4.3 and Appendix Section A.11.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See References.
    2. (b)
      Did you mention the license of the assets? [Yes] See Section A.13
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See the abstract footnote.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section A.13
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Section A.13
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Appendix
A.1 Dataset Nutrition Labels
Table 5: Definitions of the fields within the Mercury dataset.
Field Name Definition
id Task ID.
slug_name Task name.
meta_info The field accommodating the task description and submission statistics.
difficulty The difficulty level of the task.
pretty_content The field introduces the task description, examples, and constraints in pure text.
solutions Samples of solutions extracted from actual past submissions.
prompt The prompt of the solution.
entry_point The nominative entry point of the solution.
generator_code A function to generate test cases.
test_cases A collection of generated test cases.
convert_online A function to format test cases for online evaluation.
convert_offline A function to format test cases for offline evaluation.
evaluate_offline A function designed to evaluate solutions in an offline setting.
A.2 Mercury Data Distribution and Customized Data Structures
Except for all built-in Python data structures, Mercury imports another two structures to enhance the diversity and complexity as shown in Figure 4.
Figure 4: Mercury supports two customized data structures: TreeNode and ListNode.
Splits Easy Medium Hard Sum
Mercury-train 446 968 219 1,633
Mercury-eval 88 81 87 256
Table 6: Mercury-eval encompasses 256 tasks, the difficulty level of which has been balanced for model evaluation. Mercury-train comprises the remaining 1,633 tasks for model training.
A.3 Sandbox Details
Time and Memory Limitation.
Each executed code within the sandbox is subject to certain constraints to ensure fair utilization of resources and to prevent any single code from monopolizing the system resource. Specifically, there are two primary constraints: a time limit and a memory limit. The time limit restricts how long the code can execute before being forcibly terminated, thereby ensuring that no infinite loops or excessively long computations negatively impact the availability of the sandbox. The memory limit caps the amount of RAM that a process can consume. This measure precludes a single code from exhausting the memory resources, which could lead to a denial of service for subsequent codes. In our experiment settings, the timeout limit is 30 seconds, and the memory limit is 2048 MB for each solution execution.
IO Restriction.
To mitigate harmful activities such as unauthorized command execution or data exfiltration, the sandbox imposes strict Input/Output (IO) restrictions. These restrictions include limitations on reading from or writing to the disk and restrictions on the use of network sockets for sending or receiving data. By controlling the IO operations, the sandbox can prevent many common vulnerabilities and ensure that the code runs without interfering with other processes of the host system.
Isolated File System.
The sandbox employs an isolated file system to provide a safe execution environment for the code. This means that the process running in the sandbox has its virtual file system, which is separated from the host’s file system. The isolated nature of this file system ensures that even if a process within the sandbox attempts to modify or delete files, these changes will not affect the host system or other sandboxes. It acts as a security layer, protecting the host from potential threats and maintaining the integrity of the overall system.
System Libraries Redirection.
To maintain a consistent and controlled environment, the sandbox redirects calls to system libraries to sandbox-specific versions. This is done to prevent code from using certain functions directly from the host’s system libraries, which could result in unpredictable behavior or security vulnerabilities. The redirected libraries are often limited to a subset of functionalities deemed safe and necessary for executing programs within the sandbox, thus enforcing the security policies and ensuring that the running programs behave as expected.
Single-threaded Evaluation.
Single-threaded evaluation refers to executing code using a sole thread of execution, thereby simplifying resource management and timing assessments, and mitigating the intricacies linked with multi-threaded execution, such as synchronization issues, race conditions, and potential deadlocks. This mode of operation is especially important in testing environments where reproducibility and fairness are paramount, ensuring that each piece of code is evaluated using identical computational resources.
Code Efficiency Measurement.
Figure 5 shows the overview of the code execution pipeline. We gauge the Solution Instantiation and Test Ease Evaluation time spans as the execution runtime.
Figure 5: Sandbox Execution Pipeline. 1) Test Case Generation. We first employ the corresponding test case generator for each task to produce a comprehensive set of test cases for the subsequent evaluation. 2) Context Initialization. To prevent any unexpected code behavior, the sandbox environment is meticulously reinitialized for each new task. This phase ensures that all the common libraries required for executing the solution are loaded. 3) Solution Instantiation. The solution under evaluation will be encapsulated as a solution class. 4) Test Case Evaluation. Each test case the generator provides will be rigorously executed against the solution. A solution must successfully pass all the test cases to be deemed valid. 5) Clean up. The final stage involves the sandbox dutifully clearing the namespace environment and the temporary directory. Mercury records the time consumed during the stage of Solution instantiation and Test Ease Evaluation as the primary metric for assessing code efficiency.
A.4 DPO Experiment Details
Dataset Construction.
For every task problem $T^{i}$ in Mercury, we randomly selected two solutions from the task solution set $\{s^{i}_{w},s^{i}_{l}\}\sim T^{i}_{solution}$, to construct the preference dataset $D=\{P^{i},s^{i}_{w},s^{i}_{l}\}$, where $p^{i}$ is the prompt, $s^{i}_{w}$ has a faster runtime than $s^{i}_{l}$.
Model Initialization.
RLHF [48] typically begins with a reference LLM $\pi_{ref}$. Here, we initialize $\pi_{ref}$ by maximizing the likelihood of faster code completions $(p,s_{w})\sim D$, so that $\pi_{ref}=\arg\max_{\pi}E_{(p,s_{w})\sim D}\left[\log\pi(s_{w}|p)\right]$. This procedure helps mitigate the distribution shift between the true reference distribution and $\pi_{ref}$.
Optimization.
We optimize the target LLM $\pi_{\theta}$ to minimize $\mathcal{L}_{DPO}$ for the given $\pi_{ref}$ and $D$ and desired hyperparameter $\beta$. The gradient with respect to the parameters $\theta$ can be written as $\nabla_{\theta}\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})$.
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})=-E_{(x,s_{w},s_{l})\sim D}\left[\log\alpha(\beta\log\frac{\pi_{\theta}(s_{w}|p)}{\pi_{ref}(s_{w}|p)})-\log\frac{\pi_{\theta}(s_{l}|p)}{\pi_{ref}(s_{l}|p)})\right]$ (2)
$\nabla_{\theta}\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})=\\ -\beta E_{(p,s_{w},s_{l})\sim D}\left[\underbrace{\alpha(\hat{r}_{\theta}(p,s_{l})-\hat{r}_{\theta}(p,s_{w}))}_{\textit{higher weight for wrong estimate}}\left[\underbrace{\nabla_{\theta}\log\pi(s_{w}|p)}_{\textit{increase likelihood of $s_{w}$}}-\underbrace{\nabla_{\theta}\log\pi(s_{l}|p)}_{\textit{decrease likelihood of $s_{l}$}}\right]\right]$ (3)
Intuitively, the gradient of the loss function $\mathcal{L}_{DPO}$ increases the likelihood of the preferred completions $s_{w}$ and decreases the likelihood of dis-preferred completions $s_{l}$, which are weighed by how much higher the implicit reward model $\hat{r}_{\theta}$ rates the dis-preferred completions, scaled by $\beta$, i.e., how incorrectly the implicit reward model orders the completions, accounting for the strength of the KL constraint.
A.5 External Libraries Utilized in Mercury
Raw LeetCode solutions typically commence without importing shared libraries. To avoid solution failure due to absent libraries, we proactively import the libraries listed in Figure 6 during the sandbox Context Initialization phase. Note that all these libraries are imported in a temporary namespace of which the sandbox controls code behaviors.
Figure 6: External Libraries Imported in Mercury Evaluate Framework.
A.6 Model Details
Table 7: Model Details. We evaluated LLMs ranging from 1.3B to 34B.
Model Name Model Scale Link
deepseek-coder-1.3b-base 1.3B https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base
starcoder2-3b 3B https://huggingface.co/bigcode/starcoder2-3b
deepseek-coder-6.7b-base 6.7B https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base
starcoder2-7b 7B https://huggingface.co/bigcode/starcoder2-7b
CodeLlama-7b-hf 7B https://huggingface.co/codellama/CodeLlama-7b-hf
CodeQwen1.5-7B 7B https://huggingface.co/Qwen/CodeQwen1.5-7B
CodeLlama-13b-hf 13B https://huggingface.co/codellama/CodeLlama-13b-hf
starcoder2-15b 15B https://huggingface.co/bigcode/starcoder2-15b
deepseek-coder-33b-base 33B https://huggingface.co/deepseek-ai/deepseek-coder-33b-base
CodeLlama-34b-hf 34B https://huggingface.co/codellama/CodeLlama-34b-hf
A.7 A Mercury Example
Figure 7: This case is drawn from the Mercury-eval benchmark. The upper block presents the problem statement with its example, while the subsequent portion exhibits the corresponding solutions. Although all solutions are functionally correct, they exhibit significant differences in runtimes.
A.8 A HumanEval Example
Figure 8: An HumanEval example of insufficient test cases. Even though the code passed all test cases in the dashed-line box, it remains vulnerable to timeout or stack overflow when subjected to a larger input.
A.9 Prompts for Code Generation
To guarantee a fair comparison, we apply a unified one-shot prompt template for each pre-trained Code LLM. As displayed in Figure 9, the prompt template contains one shot example as well as three placeholders: <task_content>, <code_starter>, and <code_completion>.
Figure 9: Code Generation Prompts. Lines 1 to 40 are the one-shot example. In Mercury experiments, we feed the pretty_content field to the placeholder <task_content>, the prompt field to the placeholder <code_starter>, and the solution field to the placeholder <code_completion>
A.10 Hardware-agnostic Evaluation
Figure 10: Beyond scores of ‘deepseek-coder-33b’ (solid line) and ‘deepseek-coder-6.7b’ (dashed line) across varied Intel Skylake CPU configurations. The results show that Beyond can remain consistent across different hardware configurations.
A.11 Distribution of Bootstrapped Beyond Scores
Figure 11: Bootstrapped Beyond Distribution. We evaluate 3B, 7B, and 15B Starcoder2[24] models using the Mercury benchmark. Each model was executed 50 times to ensure score robustness. The y-axis in the resulting histogram represents the frequency of observations within each bin.
A.12 Dataset Metadata
The Mercury dataset is hosted on Huggingface: https://huggingface.co/datasets/Elfsong/Mercury. The Croissant Metadata can be found at https://huggingface.co/api/datasets/Elfsong/Mercury/croissant.
A.13 Legal Compliance
In this study, we have curated a comprehensive dataset by gathering publicly accessible task descriptions and archived solutions from LeetCode (https://leetcode.com/problemset/). We have ensured that our collection process is strictly limited to tasks available in the free domain, intentionally excluding any content that falls under the paid services of the platform. We abide by Fair Use [33] (Section 107): “the fair use of a copyrighted work, including such use by … scholarship, or research, is not an infringement of copyright”, where fair use is determined by “the purpose and character of the use, including whether such use is of a commercial nature or is for nonprofit educational purposes”. With the Mercury dataset, we emphasize its strictly non-commercial nature and underscore its purpose: to facilitate and advance academic research. The Mercury dataset is released under Creative Commons Attribution Non Commercial 4.0 [10].